OpenStudy (anonymous):

the 3rd term of an arithmetic sequence is 17, and 9th is 83. what is the 1st term

OpenStudy (anonymous):

This is an arithmetic progression (AP) The general formula for any term is \[a _{n} = a _{1} + (n-1)d\] a(n) is the term you want, a1 is the first term and 'd' is the difference between terms so for third term, 17 = a1+2d ===> a1 = 17-2d for 9th term 83 = a1 + 8d Solve the equation simultaneously 83= 17 - 2d +8d 6d = 66 d = 11 a1 = 17 -2(11) = -5 First term is -5